# Advanced Geometric Correction Algorithms for 3D Medical Reconstruction: Comparison of Computed Tomography and Macroscopic Imaging

### v1

@@ 4 Method of Matching Scans
The method for matching kidney sections employs an algorithmic approach to register each successive section relative to the previous one, aiming to create a continuous sequence of images that will enable three-dimensional reconstruction. This registration framework is built upon a slice-wise 2D formulation and a non-diffeomorphic deformation assumption. This scope is chosen for practical reasons, as physical macroscopic sections often exhibit discontinuous tissue distortions and tears that violate the topology-preserving constraints of diffeomorphic models. While these assumptions prioritize robustness in handling physical artifacts, the framework could be extended in future work to include 3D volumetric interpolation or diffeomorphic displacement fields (e.g., stationary velocity fields) to enhance topological consistency.
This process requires precise correction of spatial distortions resulting from tissue preparation non-uniformities.
@@
Figure 3: Symbolic illustration of the OCM algorithm’s processing steps. Beginning with the first image (1), each subsequent image is aligned through translation, rotation, and scaling transformations to produce a seamless sequence (2-5).
@@
1. Scaling: Scaling is accomplished by resizing the section. If $s<1$, the section is reduced, and for $s>1$, it is enlarged.
2. Rotation: The rotation of section $I_{i}$ by angle $\theta$ is described by the rotation matrix $R(\theta)$ on equation (3)
@@
3. Translation: Translation is performed by shifting the section’s pixels by vector $t_{x},t_{y}]$.
@@
Limitations in the Method of Matching Kidney Sections: In the process of matching kidney sections, we apply a specialized optimization technique that allows for the consideration of lower and upper bounds for transformation parameters. This method is essential for maintaining realistic transformation parameters, such as scaling, rotation, and translation.
@@
- $LB$ and $UB$: These are the lower ($LB$ - Lower Bound) and upper ($UB$ - Upper Bound) limits for our variable $x$. They define the range within which the variable $x$ can vary. For example, if the rotation can only be within the range of -45 to +45 degrees, then $LB$ would be -45, and $UB$ would be +45.
@@
Figure 4 presents a set of kidney masks in CT imaging, showing a clear and uniform structure of sections. This continuity of sections is crucial for effective three-dimensional reconstruction and allows for precise alignment of successive layers.
@@
By applying advanced image processing techniques and optimization, significant improvements in section matching were achieved, which is crucial for accurately replicating the structure of the kidneys and precise 3D reconstruction. Figures 4, 5, and 6 illustrate the evolution of the process from raw sections to the final matching outcome, highlighting the efficiency of the applied method.
The integration of this optimization technique into the section matching process allows for accurate and realistic adjustment of each kidney section, which is essential for creating precise and consistent 3D models. It also ensures control over the transformation parameters, ensuring that transformations are physically possible and consistent with the actual properties of the imaged tissues.
To better demonstrate the effectiveness of our kidney section matching method, we include visualizations of 3D kidney models before and after applying our process. The figure 7 on the left side displays a 3D kidney model obtained before applying our matching method. In this image, some discontinuity and inconsistency in the model’s structure can be noticed, resulting from suboptimal matching of individual sections.
The figure 7 on the right side demonstrates a 3D kidney model after applying our section matching process. Thanks to the use of advanced optimization techniques and precise section matching, the 3D model shows a significant improvement in uniformity and structural consistency. As a result, we obtain a model that is much more faithful to the actual anatomy of the kidney, which is crucial for medical and research applications.
@@
Two fundamentally distinct processes contribute to the dimensional accuracy of the reconstructed 3D kidney models. First, metric calibration is performed using the Hough Transform, which detects the calibration grid embedded in each macroscopic image. This grid provides a physical pixel-to-millimeter conversion ratio specific to each slice. The resulting scale factor is constant per image and is exclusively used to convert pixel-based dimensions and volumes into real-world physical units (e.g., millimeters or cubic centimeters). Importantly, this calibration step is entirely independent of any optimization procedure and is not modified throughout the reconstruction process. Second, the Optimal Cross-section Matching (OCM) algorithm applies geometric alignment transformations-namely, translation, rotation, and a local mask-based scaling-to improve spatial consistency between adjacent slices. The purpose of this internal scaling is not to modify physical units, but to enhance anatomical continuity by allowing slight morphological adjustments to the kidney contours. These adjustments are implemented using morphological operations such as imerode and imdilate, which affect the shape of the masks without altering image resolution or affecting real-world measurements. The geometric transformation parameters are constrained to realistic anatomical bounds, ensuring that local modifications remain plausible.
@@ 5 Application of the Hough Transform for Metric Grid Calibration in Macroscopic Kidney Image Analysis
The Hough Transform [undefo], utilized in our study, served to precisely detect metric grid lines on macroscopic images of kidneys. This method enabled accurate measurement of the metric grid lines’ lengths, allowing for the assessment of the kidneys’ actual sizes.
The Hough Transform is defined as a function that transforms points from the image space (xy) to a new parameter space, where each point in the image is represented by a set of curves in the parameter space. Lines in the image are identified as points of intersection of these curves, which mathematically describes the equation (7).
@@
where $\rho$ is the distance from the origin to the line, and $\theta$ is the angle of inclination of the line to the x-axis.
5.1 Peak Detection in the Hough Transform
@@
$P=\text{houghpeaks}(H,N,\text{'threshold'},\tau)$ (8)
@@ 6.1 Application of Bézier Curves
Bézier curves [undefl] were used to interpolate and smoothly connect key points that define the shape and boundaries of the kidneys and potential pathological changes within them. Through the use of these curves, it was possible to create clear and intuitive visualizations that facilitate the identification and assessment of the morphology of the examined structures.
@@
6.2 Deep Learning–Based Registration with VoxelMorph
@@
Following Balakrishnan et al. [undefd], the network predicts a dense displacement field $\phi_{i}$ to warp the pre-aligned slice $I^{\prime}_{i}=T_{OCM}(I_{i})$ to $I_{i-1}$. We use a compact 2D U-Net with the unsupervised loss:
@@ 7 Automated Kidney Reconstruction from CT and Macroscopic Image Analysis: Experimental Validation
In this chapter we present the results of experiments conducted to assess the effectiveness of our developed method for the automatic reconstruction of three-dimensional kidney models based on computed tomography (CT) images and macroscopic images, supported by appropriate organ masks. These studies are crucial for evaluating the reliability and precision of the transformation used to create 3D models, representing an innovative approach to the analysis and modeling of anatomical structures in medicine.
@@ 7.1 Results
The quantitative results for all 40 patients are summarized in Table X. The OCM+DL framework achieved a mean Dice similarity coefficient of 0.900 ($\pm$ 0.010), significantly outperforming the DL-only baseline, which yielded a mean of 0.783 ($\pm$ 0.029). The Wilcoxon signed-rank test confirmed the statistical significance of this improvement ($p<0.001$). As shown in the boxplot (Fig. 9), the OCM+DL method not only improved the overall accuracy but also substantially reduced the variability across the dataset, demonstrating higher consistency compared to the DL-only approach.
@@ 7.1.1 Evaluation Metrics
Overlap-based metrics (binary masks).
@@
Geometric agreement (binary masks).
Let principal axes $(o_{1},o_{2},o_{3})$ be obtained by PCA on the CT mask $B$. For any mask $M\in\{A,B\}$ and axis $o_{k}$, define extent:
@@
Intensity-based similarity (optional).
When comparing continuous images (not masks), we use local $\mathrm{NCC}\uparrow$ and $\mathrm{SSIM}\uparrow$ averaged over the kidney region.
@@ 7.1.3 OCM-only: Dimensional and Volumetric Validation
Table 2 illustrates L/T/W measured on transformed macroscopic images (OCM) versus CT for a representative subset of 5 patients.
@@ 7.1.5 Final Alignment: OCM + DL (Two-stage)
We next evaluate the two-stage pipeline, where OCM provides global alignment and DL performs residual local refinement between consecutive slices. We compare OCM-only, DL-only, and OCM+DL using mask-based metrics and geometric DCs (primary analysis), as well as runtime/data efficiency.
Method NCC↑ SSIM↑ Dice↑ IoU↑ HD95 (mm)↓ DC(Vol)↓
DL-only 0.85 ± 0.07 0.65 ± 0.06 0.78 ± 0.06 0.68 ± 0.04 2.9 ± 1.6 0.25 ± 0.05
@@ 8 Discussion
The proposed two-stage pipeline-combining the deterministic Optimal Cross-section Matching (OCM) and a lightweight VoxelMorph-style deep learning (DL) refinement demonstrates a practical and robust strategy for macroscopic-to-CT registration of kidney models. The integration of these two complementary approaches allows for precise geometric alignment while maintaining computational efficiency and interpretability. OCM provides global normalization by correcting scale, rotation, and translation errors, thereby establishing a consistent anatomical frame of reference. The DL-based component then refines local slice-to-slice deformations that cannot be captured by global affine transformations alone. This combination substantially reduces residual geometric discrepancies and yields smoother and more anatomically faithful 3D reconstructions.
The evaluation results confirm that the hybrid OCM+DL approach achieves the highest geometric and visual accuracy across all tested configurations. Quantitatively, it delivers higher Dice and IoU scores, lower HD95 boundary errors, and smaller dimension and volume difference coefficients (DC) compared with OCM-only and DL-only variants. These findings indicate that combining interpretable deterministic registration with learned residual refinement yields both accuracy and efficiency gains. The OCM-only method efficiently removes large-scale misalignments, while the DL module focuses on fine local corrections, leading to faster convergence, shorter inference times, and reduced training data requirements. In particular, the OCM+DL model requires significantly fewer training pairs than DL-only, since the network operates on pre-aligned input with limited residual deformation. This design aligns with the principles of residual learning and amortized optimization observed in recent registration frameworks.
@@
Overall, this study demonstrates that a synergistic combination of interpretable geometric correction (OCM) and data-driven refinement (DL) provides a powerful and generalizable framework for multimodal anatomical registration. The findings reinforce the broader concept that hybrid models-integrating deterministic and learning-based components-can effectively balance accuracy, interpretability, and efficiency in medical image analysis.
A key advantage of the OCM+DL framework is its robustness in challenging anatomical scenarios. In an analysis of the ”difficult cases” (defined as the 5 patients with the lowest baseline performance), the mean Dice score increased from 0.738 to 0.902. This average improvement of 16.4% suggests that the integration of OCM effectively compensates for limitations in standard deep learning models when dealing with high-variability cases.
@@
From a computational perspective, the hybrid framework offers a favorable trade-off between accuracy and speed, which is critical for clinical adoption in surgical planning. As detailed in Table 1, the total reconstruction time for a complete 3D kidney model is approximately 3 minutes on a standard GPU-enabled workstation. While the OCM stage remains the primary bottleneck due to its iterative nature ( 12 s per slice), the near-instantaneous inference of the DL refinement network (<<0.1 s) ensures that the pipeline remains viable for time-sensitive diagnostic workflows. This efficiency allows for rapid intraoperative or postoperative morphological assessment, bridging the gap between high-fidelity 3D modeling and the fast-paced requirements of modern surgical environments without requiring massive computational clusters.
@@
9 Conclusion
This paper presented a novel hybrid registration framework that integrates Optimal Cross-section Matching (OCM) with a VoxelMorph-style deep learning refinement to align macroscopic kidney images with CT-based anatomical references. The two-stage design leverages OCM for global alignment and deep learning for residual local correction, resulting in improved overlap (Dice, IoU), reduced boundary error (HD95), and enhanced geometric consistency (dimension and volume DC). Quantitative evaluation across 40 patients demonstrated that OCM alone effectively corrects large-scale misalignments, while the addition of DL refinement yields further gains in anatomical accuracy and spatial coherence.
The OCM+DL configuration outperformed both OCM-only and DL-only baselines, achieving the best trade-off between precision, runtime, and data efficiency. This outcome highlights the benefit of combining deterministic and learned components in a unified framework: OCM provides interpretability and stability, whereas the DL module captures subtle nonrigid deformations in a data-efficient manner. Such hybrid registration methods are particularly valuable in clinical and educational contexts where high anatomical fidelity and computational reliability are required.
@@ 9 Conclusion
The data that support the findings of this study are available from the Military Medical Institute in Warsaw, but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of the Military Medical Institute.
The data supporting the reported results are private and belong to the hospital. Due to privacy and ethical restrictions, the data cannot be shared publicly.

### v2

@@ 4 Method of Matching Scans
The method for matching kidney sections employs an algorithmic approach to register each successive section relative to the previous one, aiming to create a continuous sequence of images that will enable three-dimensional reconstruction. This registration framework is built upon a slice-wise 2D formulation and a non-diffeomorphic deformation assumption. This scope is chosen for practical reasons, as physical macroscopic sections often exhibit discontinuous tissue distortions and tears that violate the topology-preserving constraints of diffeomorphic models. While these assumptions prioritize robustness in handling physical artifacts, the framework could be extended in future work to include 3D volumetric interpolation or diffeomorphic displacement fields (e.g., stationary velocity fields) to enhance topological consistency. This process requires precise correction of spatial distortions resulting from tissue preparation non-uniformities.
@@
Figure 3: Symbolic illustration of the OCM algorithm’s processing steps. Beginning with the first image (1), each subsequent image is aligned through translation, rotation, and scaling transformations to produce a seamless sequence (2–5).
@@
  1. 1.
    Scaling: Scaling is accomplished by resizing the section. If $s<1$, the section is reduced, and for $s>1$, it is enlarged.
  2. 2.
    Rotation: The rotation of section $I_{i}$ by angle $\theta$ is described by the rotation matrix $R(\theta)$ on Equation (3)
@@
  3. 3.
    Translation: Translation is performed by shifting the section’s pixels by vector $t_{x},t_{y}]$.
@@
Limitations in the Method of Matching Kidney Sections:
In the process of matching kidney sections, we apply a specialized optimization technique that allows for the consideration of lower and upper bounds for transformation parameters. This method is essential for maintaining realistic transformation parameters, such as scaling, rotation, and translation.
@@
    LBLB and $UB$: These are the lower ($LB$-Lower Bound) and upper ($UB$-Upper Bound) limits for our variable $x$. They define the range within which the variable $x$ can vary. For example, if the rotation can only be within the range of -45 to +45 degrees, then $LB$ would be -45, and $UB$ would be +45.
@@
Figure 4 presents a set of kidney masks in CT imaging, showing a clear and uniform structure of sections. This continuity of sections is crucial for effective three-dimensional reconstruction and allows for the precise alignment of successive layers.
@@
By applying advanced image processing techniques and optimization, significant improvements in section matching were achieved, which is crucial for accurately replicating the structure of the kidneys and precise 3D reconstruction. Figures 46 illustrate the evolution of the process from raw sections to the final matching outcome, highlighting the efficiency of the applied method.
The integration of this optimization technique into the section matching process allows for the accurate and realistic adjustment of each kidney section, which is essential for creating precise and consistent 3D models. It also ensures control over the transformation parameters, ensuring that transformations are physically possible and consistent with the actual properties of the imaged tissues.
To better demonstrate the effectiveness of our kidney section matching method, we include visualizations of 3D kidney models before and after applying our process. Figure 7 on the left side displays a 3D kidney model obtained before applying our matching method. In this image, some discontinuity and inconsistency in the model’s structure can be observed, resulting from suboptimal matching of individual sections.
Figure 7 on the right side demonstrates a 3D kidney model after applying our section matching process. Thanks to the use of advanced optimization techniques and precise section matching, the 3D model shows a significant improvement in uniformity and structural consistency. As a result, we obtain a model that is much more faithful to the actual anatomy of the kidney, which is crucial for medical and research applications.
@@
Two fundamentally distinct processes contribute to the dimensional accuracy of the reconstructed 3D kidney models. First, metric calibration is performed using the Hough Transform, which detects the calibration grid embedded in each macroscopic image. This grid provides a physical pixel-to-millimeter conversion ratio specific to each slice. The resulting scale factor is constant per image and is exclusively used to convert pixel-based dimensions and volumes into real-world physical units (e.g., millimeters or cubic centimeters). Importantly, this calibration step is entirely independent of any optimization procedure and is not modified throughout the reconstruction process. Second, the Optimal Cross-section Matching (OCM) algorithm applies geometric alignment transformations—namely, translation, rotation, and a local mask-based scaling—to improve spatial consistency between adjacent slices. The purpose of this internal scaling is not to modify physical units, but to enhance anatomical continuity by allowing slight morphological adjustments to the kidney contours. These adjustments are implemented using morphological operations such as imerode and imdilate, which affect the shape of the masks without altering image resolution or affecting real-world measurements. The geometric transformation parameters are constrained to realistic anatomical bounds, ensuring that local modifications remain plausible.
@@ 5 Application of the Hough Transform for Metric Grid Calibration in Macroscopic Kidney Image Analysis
The Hough Transform [Illingworth and Kittler(1988)], utilized in our study, served to precisely detect metric grid lines on macroscopic images of kidneys. This method enabled accurate measurement of the metric grid lines’ lengths, allowing for the assessment of the kidneys’ actual sizes.
The Hough Transform is defined as a function that transforms points from the image space $(x,y)$ into a new parameter space, where each point in the image is represented by a set of curves in the parameter space. Lines in the image are identified as points of intersection of these curves, mathematically described by Equation (7).
@@
where $\rho$ is the distance from the origin to the line and $\theta$ is the angle of inclination of the line to the $x$-axis.
Peak Detection in the Hough Transform
@@
$P=\text{houghpeaks}(H,N,\text{`threshold'},\tau)$ (8)
@@ 6.1 Application of Bézier Curves
Bézier curves [Han et al.(2008)Han, Ma, and Huang] were used to interpolate and smoothly connect key points that define the shape and boundaries of the kidneys and potential pathological changes within them. Through the use of these curves, it was possible to create clear and intuitive visualizations that facilitate the identification and assessment of the morphology of the examined structures.
@@
6.2 Deep Learning-Based Registration with VoxelMorph
@@
Following Balakrishnan et al. [Balakrishnan et al.(2019)Balakrishnan, Zhao, Sabuncu, Guttag, and Dalca], the network predicts a dense displacement field $\phi_{i}$ to warp the pre-aligned slice $I^{\prime}_{i}=T_{OCM}(I_{i})$ to $I_{i-1}$. We use a compact 2D U-Net with the unsupervised loss:
@@ 7 Automated Kidney Reconstruction from CT and Macroscopic Image Analysis: Experimental Validation
In this chapter, we present the results of experiments conducted to assess the effectiveness of our developed method for the automatic reconstruction of three-dimensional kidney models based on computed tomography (CT) images and macroscopic images, supported by appropriate organ masks. These studies are crucial for evaluating the reliability and precision of the transformation used to create 3D models, representing an innovative approach to the analysis and modeling of anatomical structures in medicine.
@@ 7.1 Results
The quantitative results for all 40 patients are summarized in Table 2. The OCM + DL framework achieved a mean Dice similarity coefficient of 0.900 ($\pm$0.010), significantly outperforming the DL-only baseline, which yielded a mean of 0.783 ($\pm$0.029). The Wilcoxon signed-rank test confirmed the statistical significance of this improvement ($p<0.001$). As shown in the boxplot (Figure 9), the OCM + DL method not only improved the overall accuracy but also substantially reduced the variability across the dataset, demonstrating higher consistency compared to the DL-only approach.
@@ 7.1.1 Evaluation Metrics
Overlap-based metrics (binary masks)
@@
Geometric agreement (binary masks)
Let principal axes $(o_{1},o_{2},o_{3})$ be obtained by PCA on the CT mask $B$. For any mask $M\in\{A,B\}$ and axis $o_{k}$, define the extent as
@@
Intensity-based similarity (optional)
When comparing continuous images (not masks), we use local $\mathrm{NCC}$ and $\mathrm{SSIM}$ averaged over the kidney region.
@@ 7.1.3 OCM-Only: Dimensional and Volumetric Validation
Table 3 illustrates L/T/W measured on transformed macroscopic images (OCM) versus CT for a representative subset of five patients.
@@ 7.1.5 Final Alignment: OCM + DL (Two-Stage)
We next evaluate the two-stage pipeline, where OCM provides global alignment and DL performs residual local refinement between consecutive slices. We compare OCM-only, DLhl-only, and OCM + DL using mask-based metrics and geometric DCs (primary analysis), as well as runtime/data efficiency.
@@ 8 Discussion
The proposed two-stage pipeline combining the deterministic Optimal Cross-section Matching (OCM) and a lightweight VoxelMorph-style deep learning (DL) refinement represents a practical and robust strategy for macroscopic-to-CT registration of kidney models. The integration of these two complementary approaches allows for precise geometric alignment while maintaining computational efficiency and interpretability. OCM provides global normalization by correcting scale, rotation, and translation errors, thereby establishing a consistent anatomical frame of reference. The DL-based component then refines local slice-to-slice deformations that cannot be captured by global affine transformations alone. This combination substantially reduces residual geometric discrepancies and yields smoother and more anatomically faithful 3D reconstructions.
The evaluation results confirm that the hybrid OCM + DL approach achieves the highest geometric and visual accuracy across all tested configurations. Quantitatively, it delivers higher Dice and IoU scores, lower HD95 boundary errors, and smaller dimension and volume difference coefficients (DC) compared with the OCM-only and DL-only variants. These findings indicate that combining interpretable deterministic registration with learned residual refinement yields both accuracy and efficiency gains. The OCM-only method efficiently removes large-scale misalignments, while the DL module focuses on fine local corrections, leading to faster convergence, shorter inference times, and reduced training data requirements. In particular, the OCM + DL model requires significantly fewer training pairs than DL-only, since the network operates on pre-aligned input with limited residual deformation. This design aligns with the principles of residual learning and amortized optimization observed in recent registration frameworks.
@@
Overall, this study demonstrates that a synergistic combination of interpretable geometric correction (OCM) and data-driven refinement (DL) provides a powerful and generalizable framework for multimodal anatomical registration. The findings reinforce the broader concept that hybrid models integrating deterministic and learning-based components can effectively balance accuracy, interpretability, and efficiency in medical image analysis.
A key advantage of the OCM + DL framework is its robustness in challenging anatomical scenarios. In an analysis of the “difficult cases” (defined as the five patients with the lowest baseline performance), the mean Dice score increased from 0.738 to 0.902. This average improvement of 16.4% suggests that the integration of OCM effectively compensates for limitations in standard deep learning models when dealing with high-variability cases.
@@
From a computational perspective, the hybrid framework offers a favorable trade-off between accuracy and speed, which is critical for clinical adoption in surgical planning. As detailed in Table 1, the total reconstruction time for a complete 3D kidney model is approximately 3 min on a standard GPU-enabled workstation. While the OCM stage remains the primary bottleneck due to its iterative nature blue(about 90 s per image), the near-instantaneous inference of the DL refinement network (<<0.1 s) ensures that the pipeline remains viable for time-sensitive diagnostic workflows. This efficiency allows for rapid intraoperative or postoperative morphological assessment, bridging the gap between high-fidelity 3D modeling and the fast-paced requirements of modern surgical environments without requiring massive computational clusters.
@@
9 Conclusions
This paper presented a novel hybrid registration framework that integrates Optimal Cross-section Matching (OCM) with a VoxelMorph-style deep learning refinement to align macroscopic kidney images with CT-based anatomical references. The two-stage design leverages OCM for global alignment and deep learning for residual local correction, resulting in improved overlap (Dice, IoU), reduced boundary errors (HD95), and enhanced geometric consistency (dimension and volume DC). Quantitative evaluation across 40 patients demonstrated that OCM alone effectively corrects large-scale misalignments, while the addition of DL refinement yields further gains in anatomical accuracy and spatial coherence.
The OCM + DL configuration outperformed both the OCM-only and DL-only baselines, achieving the best trade-off between precision, runtime, and data efficiency. This outcome highlights the benefit of combining deterministic and learned components in a unified framework: OCM provides interpretability and stability, whereas the DL module captures subtle nonrigid deformations in a data-efficient manner. Such hybrid registration methods are particularly valuable in clinical and educational contexts where high anatomical fidelity and computational reliability are required.
@@ 9 Conclusions
The data
The data that support the findings of this study are available from the Military Institute of Medicine-National Research Institute in Warsaw, but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of the Military Institute of Medicine-National Research Institute.